1/2

Christian Morgenstern,  Christian Morgenstern,
The Banshee (An Approach)The Werewolf
translated by Max Knight  translated by Alexander Gross

One night, a banshee slunk away from mate and child, and in the gloom went to a village teacher’s tomb, requesting him: “Inflect me, pray.”

A Werewolf, troubled by his name, Left wife and brood one night and came To a hidden graveyard to enlist The aid of a long-dead philologist.

The village teacher climbed up straight upon his grave stone with its plate and to the apparition said who meekly knelt before the dead:

“Oh sage, wake up, please don’t berate me,”

He howled sadly, “Just conjugate me.”

The seer arose a bit unsteady

Yawned twice, wheezed once, and then was ready.

“The banSHEE, in the subject’s place; the banHERS, the possessive case. The banHER, next, is what they call objective case–and that is all.”

“Well, ‘Werewolf’ is your plural past, While ‘Waswolf’ is singularly cast: There’s ‘Amwolf’ too, the present tense, And ‘Iswolf,’ ‘Arewolf’ in this same sense.”

The banshee marveled at the cases and writhed with pleasure, making faces, but said: “You did not add, so far, the plural to the singular!”

“I know that–I’m no mental cripple–

The future form and participle

Are what I crave,” the beast replied. The scholar paused–again he tried:

The teacher, though, admitted then that this was not within his ken. “While bans are frequent”, he advised, “A she cannot be plurized.”

“A ‘Will-be-wolf?’ It’s just too long:

‘Shall-be-wolf?’ ‘Has-been-wolf?’ Utterly wrong! Such words are wounds beyond all suture– I’m sorry, but you have no future.”

The banshee, rising clammily, wailed: “What about my family?” Then, being not a learned creature, said humbly “Thanks” and left the teacher.

The Werewolf knew better–his sons still slept At home, and homewards now he crept, Happy, humble, without apology For such folly of philology.

…  …

2/2

Translation of the German poem “Der Werwolf” by Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) into English by Max Knight (civila nomo: Max Kühnel, ∗1909-06-08 – †1993-08-31).

Arg-825-1660 (2013-02-13 22:36:06)

Tiu ĉi angligo troviĝas en http: // bertilow. com/ literaturo/ lupfantomo. html . Pri la tradukinto Max Knight vidu: http: // www. jbeilharz. de/ morgenstern/ morgenstern_poems. html kaj http: // www. 123people. de/ ext/ frm? ti= ersonensuche% 20telefonbuch&search_term= m x% 20k% C3% BChnel&search_country= DE&st= suc e% 20nach% 20personen&target_url= aHR0cDovL d3dy5saXRlcmF0dXJoYXVzLmF0L2luZGV4LnBocD9pZD02ODc4 .

Translation of the German poem “Der Werwolf” by Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) into English by Alexander Gross.

Arg-825-1657 (2013-02-13 18:39:45)

Tiu ĉi angligo troviĝas en http: // bertilow. com/ literaturo/ lupfantomo. html .

http://www.poezio.net/version?poem-id=825&version-id=1660,1657